# Chain-linked Multiple Matrix Integration via Embedding Alignment

### v2

@@ 3 Theoretical Results
$\mathbf{D}^{(i)}_{k_{1},k_{2}}:=[\mathrm{Var}(\mathbf{N}^{(i)}_{k_{1},k_{2}})+(1-q_{i})(\mathbf{P}_{k_{1},k_{2}}^{(i)})^{2}]/q_{i}\text{ for any }k_{1},k_{2}\in[n_{i}].$
@@
For any fixed $(s,t)\in[n_{i_{0}}]\times[n_{i_{L}}]$, define $\tilde{\sigma}^{2}_{s,t}$ as
$\tilde{\sigma}^{2}_{s,t}:=\sum_{k_{1}=1}^{n_{i_{0}}}(\mathbf{B}^{(i_{0},i_{L})}_{k_{1},t})^{2}\mathbf{D}^{({i_{0}})}_{s,k_{1}}+\sum_{k_{2}=1}^{n_{i_{L}}}(\mathbf{B}^{(i_{L},i_{0})}_{k_{2},s})^{2}\mathbf{D}^{({i_{L}})}_{t,k_{2}}.$
@@
where $\bm{x}_{i_{0},s}$ and $\bm{x}_{i_{L},t}$ denote the $s$-th row and $t$-th row of $\mathbf{X}_{\mathcal{U}_{i_{0}}}$ and $\mathbf{X}_{\mathcal{U}_{i_{L}}}$ respectively. Note that $\tilde{\sigma}^{2}_{s,t}\lesssim\zeta_{s,t}$. Suppose the following conditions
$\tilde{\sigma}^{2}_{s,t}\gtrsim\zeta_{s,t},$ (3.9)
@@
[…]
are satisfied, where $r_{\infty}$ and $s_{\infty}$ are upper bounds for $\|\mathbf{R}^{(i_{0},i_{L})}\|_{\max}$ and $\|\mathbf{S}^{(i_{0},\dots,i_{L})}\|_{\max}$ given in Theorem 2. We then have
$\tilde{\sigma}_{s,t}^{-1}\big{(}\hat{\mathbf{P}}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i_{L}}}-\mathbf{P}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i_{L}}}\big{)}_{s,t}\rightsquigarrow\mathcal{N}(0,1)$
as $\min\{n_{i_{0}},n_{i_{L}}\}\rightarrow\infty$.
@@
If $q_{i}=1$ then all terms depending on $\|\mathbf{P}_{\mathcal{U}_{i},\mathcal{U}_{i}}\|_{\max}$ are dropped from these conditions. Specifically, $\gamma_{i}$ in Assumption 1 simplifies to $\sigma_{i}\log^{1/2}n_{i}$, and the condition in Eq. (3.10) simplifies to $\frac{\|\mathbf{X}_{\mathcal{U}_{i}}\|_{2\to\infty}\log^{1/2}n_{i}}{(n_{i}\mu_{i})^{1/2}}=o(1)$ for $i=\{i_{0},i_{L}\}$. If $\sigma_{i}=0$ then all terms depending on $\sigma_{i}$ are also dropped. Specifically, the condition in Eq. (3.10) simplifies to $\frac{\|\mathbf{X}_{\mathcal{U}_{i}}\|_{2\to\infty}\log^{1/2}(q_{i}n_{i})}{(q_{i}n_{i}\mu_{i})^{1/2}}=o(1)\text{ for }i=\{i_{0},i_{L}\}$.
@@
Using Theorem 3 we can also construct a $(1-\alpha)\times 100\%$ confidence interval for $\big{(}\mathbf{P}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i_{L}}}\big{)}_{s,t}$ as $\big{(}\hat{\mathbf{P}}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i_{L}}}\big{)}_{s,t}\pm z_{\alpha/2}\hat{\tilde{\sigma}}_{s,t}$ where $z_{\alpha/2}$ denotes the upper $\alpha/2$ quantile of the standard normal distribution and $\hat{\tilde{\sigma}}_{s,t}$ is a consistent estimate of ${\tilde{\sigma}}_{s,t}$ based on
$\hat{\mathbf{B}}^{(i_{0},i_{L})}=\hat{\mathbf{X}}_{\mathcal{U}_{i_{0}}}(\hat{\mathbf{X}}_{\mathcal{U}_{i_{0}}}^{\top}\hat{\mathbf{X}}_{\mathcal{U}_{i_{0}}})^{-1}\tilde{\mathbf{W}}^{(i_{0},i_{L})}\hat{\mathbf{X}}_{\mathcal{U}_{i_{L}}}^{\top},\quad\hat{\mathbf{B}}^{(i_{L},i_{0})}=\hat{\mathbf{X}}_{\mathcal{U}_{i_{L}}}(\hat{\mathbf{X}}_{\mathcal{U}_{i_{L}}}^{\top}\hat{\mathbf{X}}_{\mathcal{U}_{i_{L}}})^{-1}(\tilde{\mathbf{W}}^{(i_{0},i_{L})})^{\top}\hat{\mathbf{X}}_{\mathcal{U}_{i_{0}}}^{\top},$
$\hat{\mathbf{D}}^{(i_{0})}_{k_{1},k_{2}}=(\mathbf{A}^{(i_{0})}_{k_{1},k_{2}}-\hat{\mathbf{P}}^{(i_{0})}_{k_{1},k_{2}})^{2},\quad\hat{\mathbf{D}}^{(i_{L})}_{k_{1},k_{2}}=(\mathbf{A}^{(i_{L})}_{k_{1},k_{2}}-\hat{\mathbf{P}}^{(i_{L})}_{k_{1},k_{2}})^{2}$
with $\hat{\mathbf{P}}^{(i_{0})}=\hat{\mathbf{X}}^{(i_{0})}\hat{\mathbf{X}}^{(i_{0})\top},\hat{\mathbf{P}}^{(i_{L})}=\hat{\mathbf{X}}^{(i_{L})}\hat{\mathbf{X}}^{(i_{L})\top}$ and $\tilde{\mathbf{W}}^{(i_{0},i_{L})}=\mathbf{W}^{(i_{0},i_{1})}\mathbf{W}^{(i_{1},i_{2})}\dots\mathbf{W}^{(i_{L-1},i_{L})}$; we leave the details to the interested readers.
@@
We now provide an example to illustrate the above results. We first assume $\{n_{i}\}$ are of the same order, i.e., there exists an $n$ with $n_{i}\asymp n$ for all $i$. We also assume $q_{i}\equiv q$, $\sigma_{i}\equiv\sigma$ for all $i$. We further suppose $\mathbf{P}\in\mathbb{R}^{N\times N}$ has $\Theta(N^{2})$ entries that are lower bounded by some constant $c_{0}$ not depending on $N$. Then, as $\mathbf{P}$ is also low-rank with bounded condition number, we have $\lambda_{\min}=\Theta(N)$, and thus $\|\mathbf{X}\|_{2\to\infty}\lesssim 1,\|\mathbf{P}\|_{\max}\lesssim 1$. By Eq. (3.3) we have $\lambda_{i,\min}\asymp\frac{n_{i}}{N}\lambda_{\min}$, so $\mu_{i}\asymp 1$ for all $i$. For the overlaps we assume $n_{i_{\ell-1},i_{\ell}}\equiv m\geq d$ and $\vartheta_{i_{\ell-1},i_{\ell}}\asymp\theta_{i_{\ell-1},i_{\ell}}=\Theta(m)$. Under this setting, the condition in Eq (3.2) simplifies to
$\frac{(1+\sigma)\log^{1/2}n}{(qn)^{1/2}}\ll 1;$
the error bounds in Eq. (3.8) of Theorem 2 simplify to
@@
with high probability. Furthermore we also have
$\|\hat{\mathbf{P}}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i_{L}}}-\mathbf{P}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i_{L}}}\|_{\max}\lesssim\frac{(1+\sigma)\log^{1/2}n}{(qn)^{1/2}}$
with high probability, provided that
$L\Big{(}\frac{(1+\sigma)\log^{1/2}n}{(qn)^{1/2}}+\frac{1}{m^{1/2}}\Big{)}\lesssim 1.$
All conditions are then trivially satisfied and the estimate error converges to 00 whenever $\sigma=O(1)$, $L=O(1)$, and $qn=\Omega(\log^{2}n)$. Note that the overlap size $m$ can be as small as the minimal $d$.
For the normal approximation in Theorem 3, the condition in Eq. (3.10) is trivial, and the condition in Eq. (3.11) simplifies to
@@
Note that $\frac{(1+\sigma)}{(\sigma^{2}+(1-q))^{1/2}}$ is bounded when $q\leq 1-c$ for some constant $c>0$ (this condition is dropped when $q=1$). Then Eq. (3.13) is satisfied when $qn=\omega(\log^{2}n)$ and $m=\omega(\log n)$, i.e. the overlap size is slightly larger than the minimal overlap.
3.1 Comparison with related work
As mentioned in Section 2, our theoretical results are comparable to Zhou et al. (2023) for two observed submatrices and to Bishop and Yu (2014) for a simple chain of observed submatrices. In particular, while the error bounds in Zhou et al. (2023) are in the spectral norm and those in Bishop and Yu (2014) are in the Frobenius norm, our error bounds are in the maximum entrywise norm and allow for heterogeneity among blocks. In additional, the error bound in Theorem 4 of Bishop and Yu (2014) grows exponentially with the length of the chain (due to its dependency on $G^{k-2}$ where $k$ is the chain length), whereas our error bound for $\hat{\mathbf{P}}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i_{L}}}$ includes only a non-dominant term that grows linearly with the chain length; see Eq. (3.8) or Eq. (3.12). And the bound in Theorem 4 of Bishop and Yu (2014) is only applicable for small $\epsilon$ where $\|\mathbf{A}^{(i)}-\mathbf{P}^{(i)}\|_{F}\leq\epsilon$, whereas our noise model allows $\epsilon$ to be of order $n_{i}$ with $\|\mathbf{X}\|_{F}$ of order $\sqrt{N}$, rendering their result ineffective.
@@ 3.1 Comparison with related work
We emphasize that the block-wise observation models in this paper, BONMI (Zhou et al., 2023) and SPSMC (Bishop and Yu, 2014) differ significantly from those in the standard matrix completion literature, which typically focuses on a single large matrix and assumes uniformly or independently sampled observed entries. Nevertheless, the authors of BONMI have compared their results with other results in the standard matrix completion literature. For example, Remark 10 in Zhou et al. (2023) shows that the upper bound for the spectral norm error of BONMI matches the minimax rate for the missing at random setting. As CMMI is an extension of BONMI to more than $2$ matrices, the above comparison is still valid. Furthermore, our results for CMMI are in terms of the maximum entrywise norm and normal approximations, which are significant refinements of the spectral norm error in BONMI, and are thus also comparable with the best available results for standard matrix completion such as those in Abbe et al. (2020) and Chen et al. (2021). More specifically, consider the case of $n_{i}\equiv n\asymp N/L$ with $\lambda_{i,\min}\asymp\frac{n}{N}\lambda_{\min}\asymp\frac{1}{L}\lambda_{\min}$. Also suppose $q_{i}\equiv q$ and $\sigma_{i}\equiv\sigma$. Then CMMI has the maximum entrywise error bound of $\frac{(\|\mathbf{P}\|_{\max}+\sigma)\log^{1/2}N}{q^{1/2}(N/L)^{1/2}\lambda_{\min}^{1/2}}$, which matches the rate in Theorem 3.4 of Abbe et al. (2020) and Theorem 4.5 of Chen et al. (2021) up to a factor of $L^{-1/2}$, as the number of observed entries in our model is only $1/L$ times that for the standard matrix completion models. Finally the normal approximation result in Theorem 3 is analogous to Theorem 4.12 in Chen et al. (2021), with the main difference being the expression for the normalizing variance as our model considers individual noise matrices $\mathbf{E}^{(i_{0})}$ and $\mathbf{E}^{(i_{L})}$ whereas Chen et al. (2021) consider a global noise matrix $\mathbf{E}$ (which includes $\mathbf{E}^{(i_{0})}$ and $\mathbf{E}^{(i_{L})}$ as submatrices).
@@
Another related work is Chang et al. (2022) which considers matrix completion for sample covariance matrices with a spiked covariance structure. Sample covariance matrices differ somewhat from the data matrices considered in our paper as, while both our population data matrix $\mathbf{P}$ in Section 2 and their population covariance matrix $\bm{\Sigma}$ are positive semidefinite, the entries of the sample covariance matrix $\hat{\bm{\Sigma}}$ are dependent. Consequently, the settings in the two papers are related but not directly comparable. Nevertheless, if we were to compare our results against Theorem C.1 in Chang et al. (2022) (where we set $q_{i}\equiv 1$ in our model, as Chang et al. (2022) assume that the sample covariance submatrices are observed completely) then (1) we allow block sizes $\{p_{k}\}$ (theirs $\{p_{k}\}$ are our $\{n_{i}\}$) to differ significantly in magnitude; (2) more importantly, our error bounds depend at most linearly on the chain length, whereas their bounds grow exponentially with the chain length due to the dependency on $\xi^{K}$, (their $K$ is our $L$). As $\xi=\sqrt{\log\max_{i}n_{i}}$ (see Proposition C.2 in Chang et al. (2022)), this results in a factor of $(\log\max_{i}n_{i})^{K/2}$ that is highly undesirable as $K$ increases.

### v4

@@ 3 Theoretical Results
$\mathbf{D}^{(i)}_{k_{1},k_{2}}=[\mathrm{Var}(\mathbf{N}^{(i)}_{k_{1},k_{2}})+(1-q_{i})(\mathbf{P}_{k_{1},k_{2}}^{(i)})^{2}]/q_{i}\text{ for any }k_{1},k_{2}\in[n_{i}].$
@@
For any fixed $(s,t)\in[n_{i_{0}}]\times[n_{i_{L}}]$, define $\widetilde{\sigma}^{2}_{s,t}$ as
$\widetilde{\sigma}^{2}_{s,t}:=\sum_{k_{1}=1}^{n_{i_{0}}}(\mathbf{B}^{(i_{0},i_{L})}_{k_{1},t})^{2}\mathbf{D}^{({i_{0}})}_{s,k_{1}}+\sum_{k_{2}=1}^{n_{i_{L}}}(\mathbf{B}^{(i_{L},i_{0})}_{k_{2},s})^{2}\mathbf{D}^{({i_{L}})}_{t,k_{2}}$.
@@
where $\bm{x}_{i_{0},s}$ and $\bm{x}_{i_{L},t}$ denote the $s$-th row and $t$-th row of $\mathbf{X}_{\mathcal{U}_{i_{0}}}$ and $\mathbf{X}_{\mathcal{U}_{i_{L}}}$ respectively. Note that $\widetilde{\sigma}^{2}_{s,t}\lesssim\zeta_{s,t}$. Suppose the following conditions
$\widetilde{\sigma}^{2}_{s,t}\gtrsim\zeta_{s,t},$ (3.12)
@@
$\begin{aligned} (r_{\infty}+s_{\infty})/\zeta_{s,t}^{1/2}=o(1)\end{aligned}$ (3.14)
are satisfied, where $r_{\infty}$ and $s_{\infty}$ are upper bounds for $\|\mathbf{R}^{(i_{0},i_{L})}\|_{\max}$ and $\|\mathbf{S}^{(i_{0},\dots,i_{L})}\|_{\max}$ given in Theorem 2. We then have $\widetilde{\sigma}_{s,t}^{-1}\big{(}\widehat{\mathbf{P}}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i_{L}}}-\mathbf{P}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i_{L}}}\big{)}_{s,t}\rightsquigarrow\mathcal{N}(0,1)$ as $\min\{n_{i_{0}},n_{i_{L}}\}\rightarrow\infty$.
@@
Eq. (3.12) provides a lower bound for the entrywise variance $\widetilde{\sigma}^{2}_{s,t}$ and is necessary as our normal approximations are for $\widetilde{\sigma}_{s,t}^{-1}(\widehat{\mathbf{P}}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i_{L}}}-\mathbf{P}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i_{L}}})_{s,t}$. Eq. (3.13) ensures that each independent component of the dominant terms for $\big{(}\widehat{\mathbf{P}}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i_{L}}}-\mathbf{P}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i_{L}}}\big{)}_{s,t}$ is not too large compared to $\widetilde{\sigma}_{s,t}$. Eq. (3.14) guarantees that the remainder terms $\|\mathbf{R}^{(i_{0},i_{L})}\|_{\max}$ and $\|\mathbf{S}^{(i_{0},\dots,i_{L})}\|_{\max}$ are negligible (when scaled by $\widetilde{\sigma}_{s,t}^{-1})$. These conditions are very mild; see Remark 5. In addition, if $q_{i}=1$ then all terms depending on $\|\mathbf{P}_{\mathcal{U}_{i},\mathcal{U}_{i}}\|_{\max}$ are dropped from these conditions. Specifically, $\gamma_{i}$ in Assumption 1 simplifies to $\sigma_{i}\log^{1/2}n_{i}$, and Eq. (3.13) simplifies to $\frac{\|\mathbf{X}_{\mathcal{U}_{i}}\|_{2\to\infty}\log^{1/2}n_{i}}{(n_{i}\mu_{i})^{1/2}}=o(1)$ for $i=\{i_{0},i_{L}\}$. If $\sigma_{i}=0$ then all terms depending on $\sigma_{i}$ are dropped. Specifically, Eq. (3.13) simplifies to $\frac{\|\mathbf{X}_{\mathcal{U}_{i}}\|_{2\to\infty}\log^{1/2}(q_{i}n_{i})}{(q_{i}n_{i}\mu_{i})^{1/2}}=o(1)\text{ for }i=\{i_{0},i_{L}\}$.
@@
Using Theorem 3 we can also construct a $(1-\alpha)\times 100\%$ confidence interval for $\big{(}\mathbf{P}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i_{L}}}\big{)}_{s,t}$ as $\big{(}\widehat{\mathbf{P}}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i_{L}}}\big{)}_{s,t}\pm z_{\alpha/2}\widehat{\widetilde{\sigma}}_{s,t}$ where $z_{\alpha/2}$ denotes the upper $\alpha/2$ quantile of the standard normal distribution and $\widehat{\widetilde{\sigma}}_{s,t}$ is a consistent estimate of ${\widetilde{\sigma}}_{s,t}$ based on
$\widehat{\mathbf{B}}^{(i_{0},i_{L})}=\widehat{\mathbf{X}}^{(i_{0})}(\widehat{\mathbf{X}}^{(i_{0})\top}\widehat{\mathbf{X}}^{(i_{0})})^{-1}\widetilde{\mathbf{W}}^{(i_{0},i_{L})}\widehat{\mathbf{X}}^{(i_{L})\top},\quad\widehat{\mathbf{B}}^{(i_{L},i_{0})}=\widehat{\mathbf{X}}^{(i_{L})}(\widehat{\mathbf{X}}^{(i_{L})\top}\widehat{\mathbf{X}}^{(i_{L})})^{-1}\widetilde{\mathbf{W}}^{(i_{0},i_{L})\top}\widehat{\mathbf{X}}^{(i_{0})\top},$
$\widehat{\mathbf{D}}^{(i_{0})}_{k_{1},k_{2}}=(\mathbf{A}^{(i_{0})}_{k_{1},k_{2}}-\widehat{\mathbf{P}}^{(i_{0})}_{k_{1},k_{2}})^{2},\quad\widehat{\mathbf{D}}^{(i_{L})}_{k_{1},k_{2}}=(\mathbf{A}^{(i_{L})}_{k_{1},k_{2}}-\widehat{\mathbf{P}}^{(i_{L})}_{k_{1},k_{2}})^{2}$
with $\widehat{\mathbf{P}}^{(i_{0})}=\widehat{\mathbf{X}}^{(i_{0})}\widehat{\mathbf{X}}^{(i_{0})\top},\widehat{\mathbf{P}}^{(i_{L})}=\widehat{\mathbf{X}}^{(i_{L})}\widehat{\mathbf{X}}^{(i_{L})\top}$ and $\widetilde{\mathbf{W}}^{(i_{0},i_{L})}=\mathbf{W}^{(i_{0},i_{1})}\mathbf{W}^{(i_{1},i_{2})}\dots\mathbf{W}^{(i_{L-1},i_{L})}$; we leave the details to the interested readers.
@@
We now provide an example to illustrate the above results. We first assume $\{n_{i}\}$ are of the same order, i.e., there exists an $n$ with $n_{i}\asymp n$ for all $i$. We also assume $q_{i}\equiv q$, $\sigma_{i}\equiv\sigma$ for all $i$. We further suppose $\mathbf{P}\in\mathbb{R}^{N\times N}$ has $\Theta(N^{2})$ entries that are lower bounded by some constant $c_{0}$ not depending on $N$, and that $\|\mathbf{P}\|_{\max}\leq c_{1}$ for some other constant $c_{1}$ not depending on $N$. Then $\|\mathbf{X}\|_{2\to\infty}\lesssim 1$, and as $\mathbf{P}$ is low-rank with bounded condition number, we also have $\lambda_{\min}=\Theta(N)$. By Eq. (3.3) we have $\lambda_{i,\min}\asymp\frac{n_{i}}{N}\lambda_{\min}$, so $\mu_{i}\asymp 1$ for all $i$. For the overlaps we assume $n_{i_{\ell-1},i_{\ell}}\equiv m\geq d$ and $\vartheta_{i_{\ell-1},i_{\ell}}\asymp\theta_{i_{\ell-1},i_{\ell}}=\Theta(m)$. Under this setting, the condition in Eq (3.2) simplifies to $\frac{(1+\sigma)\log^{1/2}n}{(qn)^{1/2}}\ll 1;$ the error bounds in Eq. (3.9) and Eq. (3.10) of Theorem 2 simplify to
@@
with high probability. Furthermore provided the condition in Eq. (3.11), we also have
$\|\widehat{\mathbf{P}}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i_{L}}}-\mathbf{P}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i_{L}}}\|_{\max}\lesssim\frac{(1+\sigma)\log^{1/2}n}{(qn)^{1/2}}$
with high probability, and Eq. (3.11) simplifies to $L\Big{(}\frac{(1+\sigma)\log^{1/2}n}{(qn)^{1/2}}+\frac{1}{m^{1/2}}\Big{)}\lesssim 1$ now.
All conditions are then trivially satisfied and the estimate error converges to 0 whenever $\sigma=O(1)$, $L=O(1)$, and $qn=\Omega(\log^{2}n)$. Note that the overlap size $m$ can be as small as $d=\mathrm{rk}(\mathbf{P})$; see Section A.2 for simulation results. For Theorem 3, the condition in Eq. (3.13) is trivial, and the condition in Eq. (3.14) simplifies to
@@
Note that $\frac{(1+\sigma)}{(\sigma^{2}+(1-q))^{1/2}}$ is bounded when $q\leq 1-c$ for some constant $c>0$ (this condition can be omitted when $q=1$). Then Eq. (3.16) is satisfied when $qn=\omega(\log^{2}n)$ and $m=\omega(\log n)$, i.e. the overlap size is slightly larger than the minimal $d=\mathrm{rk}(\mathbf{P})$. See Section E.9 for further discussion on Assumption 1 and the conditions in Theorems 1, 2, and 3.
3.1 Discussion and comparison with related work
Our theoretical results are comparable to those of Zhou et al. (2023) for two observed submatrices and to Bishop and Yu (2014) for a simple chain, while being significantly stronger than both. In particular, the error bounds in Zhou et al. (2023) and Bishop and Yu (2014) are given in terms of the spectral and Frobenius norms, which only provide coarse control over individual entries. In other words, the estimation error for each entry can only be bounded indirectly through the overall matrix norm and does not yield sharp entrywise guarantees. In contrast, our analysis leverages a $2\to\infty$ norm bound on $\widehat{\mathbf{X}}^{(i)}$, which controls row-wise fluctuations and leads directly to a bound on the maximum entrywise error of the recovered matrix. Furthermore, our analysis also reveals the dominant error terms in our estimation from which we can obtain entrywise normal approximations. This type of inferential guarantees can not be achieved using either the spectral or Frobenius norm bounds. In addition, our theoretical results allow for heterogeneity across blocks as we can handle cases where $(n_{i},\sigma_{i},q_{i})$ have different magnitudes and the overlap sizes $n_{i,j}$ can vary across submatrices. In contrast, Zhou et al. (2023) assumes all $\{n_{i}\}$ are of the same order and all $\{n_{i,j}\}$ are of the same order, and their error bound for recovering the entries related to blocks $i$ and $j$ depends on $\sigma=\max_{k\in[m]}\sigma_{k}$ rather than the specific $(\sigma_{i},\sigma_{j})$. Our analysis is thus more general and also yields sharper results under heterogeneous settings. Finally we relax the overlap requirement, as Zhou et al. (2023) requires overlaps to grow with submatrices, while we show that the overlap size can be $d$. Comparing our results to Bishop and Yu (2014), we note that the error bound in Theorem 4 of Bishop and Yu (2014) grows exponentially with the length of the chain (due to its dependency on $G^{k-2}$ where $k$ is the chain length), whereas our bound for $\widehat{\mathbf{P}}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i_{L}}}$ includes only a non-dominant term that grows linearly with the chain length; see Eq. (3.10) or Eq. (3.15). And the bound in Theorem 4 of Bishop and Yu (2014) is only applicable for small $\epsilon$ where $\|\mathbf{A}^{(i)}-\mathbf{P}^{(i)}\|_{F}\leq\epsilon$, whereas our noise model allows $\epsilon$ to be of order $n_{i}$ with $\|\mathbf{X}\|_{F}$ of order $\sqrt{N}$, rendering their result ineffective.
@@ 3.1 Discussion and comparison with related work
We emphasize that the block-wise observation models in this paper, BONMI (Zhou et al., 2023) and SPSMC (Bishop and Yu, 2014) differ significantly from those in the standard matrix completion literature, which typically focuses on a single large matrix and assumes uniformly or independently sampled observed entries. Nevertheless, the authors of BONMI have compared their results with other results in the standard matrix completion literature. For example, Remark 10 in Zhou et al. (2023) shows that the upper bound for the spectral norm error of BONMI matches the minimax rate for the missing at random setting. As CMMI is an extension of BONMI to more than $2$ matrices, the above comparison is still valid. Furthermore, our results for CMMI are in terms of the maximum entrywise norm and normal approximations, which are significant refinements of the spectral norm error in BONMI, and are also comparable to the best available results for standard matrix completion such as those in Abbe et al. (2020) and Chen et al. (2021). More specifically, consider the case of $n_{i}\equiv n\asymp N/L$ with $\lambda_{i,\min}\asymp\frac{n}{N}\lambda_{\min}\asymp\frac{1}{L}\lambda_{\min}$. Also suppose $q_{i}\equiv q$ and $\sigma_{i}\equiv\sigma$. Then CMMI has the maximum entrywise error bound of $\frac{(\|\mathbf{P}\|_{\max}+\sigma)\log^{1/2}N}{q^{1/2}(N/L)^{1/2}}$, which matches the rate in Theorem 3.4 of Abbe et al. (2020) and Theorem 4.5 of Chen et al. (2021) up to a factor of $L^{-1/2}$, as the number of observed entries in our model is only $1/L$ times that for the standard matrix completion models. Finally the normal approximation result in Theorem 3 is analogous to Theorem 4.12 in Chen et al. (2021), with the main difference being the expression for the normalizing variance as our model considers individual noise matrices $\mathbf{E}^{(i_{0})}$ and $\mathbf{E}^{(i_{L})}$ whereas Chen et al. (2021) consider a global noise matrix $\mathbf{E}$.
@@
Another related work is Chang et al. (2022) which considers matrix completion for sample covariance matrices with a spiked covariance structure. Sample covariance matrices differ somewhat from the data matrices considered in our paper as, while both our population data matrix $\mathbf{P}$ in Section 2 and their population covariance matrix $\bm{\Sigma}$ are positive semidefinite, the entries of the sample covariance matrix $\widehat{\bm{\Sigma}}$ are dependent. Consequently, the settings in the two papers are related but not directly comparable. Nevertheless, if we were to compare our results against Theorem C.1 in Chang et al. (2022) (where we set $q_{i}\equiv 1$ in our model, as Chang et al. (2022) assume that the sample covariance submatrices are observed completely) then (1) we allow block sizes $\{p_{k}\}$ (theirs $\{p_{k}\}$ are our $\{n_{i}\}$) to differ significantly in magnitude; (2) more importantly, our error bounds depend at most linearly on the chain length, whereas their bounds grow exponentially with the chain length due to the dependency on $\xi^{K}$, (their $K$ is our $L$). As $\xi=\sqrt{\log\max_{i}n_{i}}$ (see Proposition C.2 in Chang et al. (2022)), this results in a factor of $(\log\max_{i}n_{i})^{K/2}$ that is highly undesirable as $K$ increases.
We finally note that for each observed submatrix with potentially independent random missing entries, estimating the latent positions $\widehat{\mathbf{X}}^{(i)}$ using the scaled leading eigenvectors of $\mathbf{A}^{(i)}$ (as done in this paper) is rate-optimal. More specifically, from Lemma D.1 we have $\|\widehat{\mathbf{P}}^{(i)}-\mathbf{P}^{(i)}\|_{\max}\lesssim\frac{(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})\log^{1/2}n_{i}}{q_{i}^{1/2}n_{i}^{1/2}}$ with high probability, where $\widehat{\mathbf{P}}^{(i)}=\widehat{\mathbf{X}}^{(i)}\widehat{\mathbf{X}}^{(i)\top}$. This matches the best available entrywise error rates for matrix completion established in the literature, such as those in Abbe et al. (2020) and Chen et al. (2021). Lemma D.1 also implies $\|\widehat{\mathbf{P}}^{(i)}-\mathbf{P}^{(i)}\|_{F}\lesssim\frac{(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})n_{i}^{1/2}\log^{1/2}n_{i}}{q_{i}^{1/2}}$ with high probability, and is the same (up to logarithmic factor) as the oracle bound from (Candes and Plan, 2010, Eq. (III.13)). The logarithmic factor is negligible and is due mainly to the fact that Lemma D.1 yields a concentration bound for $\|\widehat{\mathbf{P}}^{(i)}-\mathbf{P}^{(i)}\|_{F}$ while Candes and Plan (2010) is for $\mathbb{E}[\|\widehat{\mathbf{P}}^{(i)}-\mathbf{P}^{(i)}\|_{F}]$. See Section A.3 of the supplementary material for empirical comparisons between the SVD-based algorithm and other matrix completion methods for estimating $\{\mathbf{X}^{(i)}\}$. Accurate initialization of $\{\widehat{\mathbf{X}}^{(i)}\}$ is crucial for the subsequent joint integration, as it leads to more precise estimates of the transformation matrices $\{\mathbf{W}^{(i,j)}\}$ and thereby contributing to improved overall recovery.